OpenStudy (anonymous):

PLEASE HELP!! Using the properties of exponents, simplify the expression. (x^4/5)/(x^2/5)

OpenStudy (mathstudent55):

\( \dfrac{a^m}{a^n} = a^{m - n} \) When you divide powers with the same base, write the same base and subtract the exponents.

OpenStudy (anonymous):

Ok I got that one! would a^5/4/a^-1/4= a^5/4-(-1/4)? +a^6/4..Or a^3/2

OpenStudy (mathstudent55):